OpenStudy (anonymous):

how do you find the arcsin of -sqrt(3)/2

OpenStudy (anonymous):

think of arcsin as an inverse function f^(-1) (x) = - sqrt(3)/2 (I would use the math editor but it doesn't display the - sign anymore) what do we know about the relationship between functions and their inverses? if f is a function and f^(-1) is the inverse then f(f^(-1))(x) = x and we know that arcsin is the inverse of sin so the question can be restated as f(f^(-1))(x) = x or f(-sqrt(3)/2) = x or, since f(x) = sin(x) in our case sin(-sqrt(3)/2) = x

OpenStudy (anonymous):

so how do you find the answer?

OpenStudy (anonymous):

in other words arcsin (-sqrt(3)/2) = x is the same thing as saying sin(x) = -sqrt(3)/2

OpenStudy (anonymous):

just use knowledge of trigonometry sin(x) = -sqrt(3)/2 (-60 degrees) using the unit circle as our guide this corresponds to a 30-60-90 triangle the x component is positive, the y component is negative http://lscc.edu/faculty/dannyt/SiteAssets/SitePages/Trigonometry%20-%20Summer%20B/UnitCircle.gif this translates to 5 pi / 3 radians or -pi/3 radian